The distinct lines y=mx+b and y=bx+m intersect at the point (m-b, 9). The sum of x intercepts of the lines is
  1. 9
  2. -41/20
  3. 41/20
  4. Insufficient data

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
mx + b= bx + m
(m-b) x = m-b
x = 1
y = m+b
Now
m-b = 1 
m+b = 9
Henve
m = 5, b = 4
Now
X intercept for 1st line = -b/m = -4/5
X intercept for 2nd line =  -m/b = -5/4 
Now add them 
-41/20
Hence option B is correct